A stock solution of Na_2S_2O_3 is prepared by addi
Answer #1

Molecular weight of Na2S2O3= 158

moles of Na2S2O3 in 20.4 gms =20.4/158=0.129 moles

molarity = moles of Na2S2O3 in liter of solution =0.129/0.25= 0.516

moles of Na2S2O3 in 25 ml of sample taken= 0.516*25/1000=0.0129

Molarity of final solution = 0.0129/0.5 =0.0258 Mol/L